# Effect of primary scalar hair on black hole’s strong lensing in Beyond Horndeski Gravity

### v1

@@ I Introduction
Horndeski theory represents the most general class of scalar-tensor theories that maintain second-order field equations while avoiding Ostrogradski instabilities [30]. Later advancements demonstrated that scalar-tensor theories could accommodate higher-order field equations without instability by utilizing degenerate Lagrangians. This led to the development of Beyond Horndeski and Degenerate Higher-Order Scalar-Tensor theories [31, 32, 33, 34, 35, 36].
Obtaining black hole solutions in these modified theories is challenging due to extended no-hair theorems, which constrain the existence of non-trivial scalar field configurations [37, 38, 39]. The no-hair conjecture posits that a black hole is entirely described by its mass $M$, angular momentum $J$, and electric charge $Q$, measurable at asymptotic infinity [40]. A black hole with an additional global charge is said to possess primary hair, whereas secondary hair implies that the metric remains fully determined by $M$, $J$, and $Q$ despite the presence of additional fields. Both cases are often categorized under the broader concept of “hairy black holes” [41, 42, 43, 44].
@@
Building on the Virbhadra-Ellis lens equation [50, 51], Bozza extended analytical lensing methods to a wide class of static and spherically symmetric spacetimes, demonstrating that logarithmic divergence of the deflection angle at the photon sphere is a common feature in such backgrounds. He also identified the primary and secondary images and reported the existence of two infinite sets of relativistic images [52]. These findings were later extended by considering finite observer and source positions [53]. In cases where the black hole is illuminated by a thin accretion disk, Tsupko applied Bozza’s method to determine the shape of higher-order images using both analytical and numerical techniques [54]. More recently, Ref. [55] demonstrated that the spacing between higher-order images (photon rings) can reveal key properties of the underlying spacetime.
@@
Strong gravitational lensing remains a powerful tool for probing black hole spacetimes in both general relativity and alternative theories of gravity. Numerous studies have explored this phenomenon in various contexts [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77].
@@
Motivated by the considerations above regarding the choice of the gravitational theory, we structure this paper as follows: In Sect. II, we provide a concise review of the theory under investigation, introducing the static, spherically symmetric, asymptotically flat black hole solution proposed within this framework and discussing its key properties. In Sect. III, we initiate our analysis by formulating the deflection angle of light in the strong lensing regime, considering both the source and the observer at finite distances from the black hole. The relevant integrals are then evaluated numerically, and we examine how the deflection angle varies with the impact parameter of the passing light rays, as perceived by the observer. Sect. IV is dedicated to the lens equation governing relativistic images of different orders. We utilize this equation to explore the response of Einstein rings to variations in the black hole’s scalar hair. Additionally, we introduce three key observables in strong gravitational lensing and employ the angular diameter of relativistic images as a tool to assess the black hole’s properties in light of the EHT data for M87* and Sgr A*. In Sect. V, we adopt an alternative approach to constraining the black hole parameters by considering a thin accretion disk as the source of illumination. We derive analytical expressions for the shapes of images of different orders and establish a connection between the observed black hole shadow and the primary image of the accretion disk. Using this framework, we impose new constraints on the black hole’s scalar hair. Furthermore, we demonstrate why higher-order images play a less significant role in astrophysical observational tests based on EHT data. Finally, we summarize our findings in Sect. VI.
@@
Throughout this study, we employ geometric units where $18\pi G=\hbar=\ell_{p}=c=$ 1, adopt the sign convention (,+,+,+)(-,+,+,+)( - , + , + , + ), and use primes to denote derivatives with respect to the radial coordinate.
@@ II The black hole solution with primary scalar hair
This section outlines the fundamental properties of the static, spherically symmetric black hole solution introduced in Ref. [45]. In the framework of Beyond Horndeski theories [78], where the shift symmetry $\tilde{\chi}\rightarrow\tilde{\chi}+\text{const.}$. and parity symmetry $\tilde{\chi}\rightarrow{-\tilde{\chi}}$ are preserved, the theory is characterized by three arbitrary functions, $G_{2}$, $G_{4}$, and $F_{4}$, which are dependent on the kinetic term of the scalar field, given by $X=-\frac{1}{2}\tilde{\chi}_{,\mu}\tilde{\chi}^{,\mu}$. The corresponding action takes the form
$I\left[g_{\mu\nu},\tilde{\chi}\right]=\frac{1}{2}\int\mathrm{d}^{4}x\,\sqrt{-g}\Biggl{\{}G_{2}(X)+G_{4}(X)R+G_{4X}\left[(\square\tilde{\chi})^{2}-\tilde{\chi}_{\mu\nu}\tilde{\chi}^{\mu\nu}\right]+F_{4}(X)\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma}{}_{\sigma}\tilde{\chi}_{\mu}\tilde{\chi}_{\alpha}\tilde{\chi}_{\nu\beta}\tilde{\chi}_{\rho\gamma}\Biggr{\}},$ (1)
@@
Since the theory remains unchanged under the transformation $\lambda\rightarrow-\lambda$, we assume $\lambda$ to be positive. Setting $M=\mathcal{c}=0$ leads to a Minkowski spacetime, whereas for $M=0$ but $\mathcal{c}\neq 0$, the solution describes a nontrivial spacetime. When $\mathcal{c}<0$, it corresponds to a black hole, while for $\mathcal{c}>0$, it gives rise to a naked singularity.
@@ III Strong gravitational lensing by the PHBH
where $-1\leq\epsilon<\infty$. Since strong lensing effects are significant for $b\approx b_{c}$ (or equivalently, $R\approx r_{p}$), corresponding to trajectories near the photon sphere, the parameter $\epsilon$ is expected to be small. In other words,
@@
Note that, as $R$ approaches $r_{p}$, then $\mathcal{R}$ approaches 00, and hence, the integrals in Eq. (12) diverge. In Refs. [52, 53], a method was developed to address this divergence using logarithmic approximations. Under these approximations, and assuming the source and observer are located at finite positions, the azimuthal angle shift is given by [55]
@@
×$\times\Biggl{\{}\sqrt{2}f_{0}\eta\Biggl{[}r_{p}\Biggl{(}-12M\left(\lambda^{2}+r_{p}^{2}\right)^{3}+6r_{p}^{7}+18\lambda^{2}r_{p}^{5}+18\lambda^{4}r_{p}^{3}+6\lambda^{6}r_{p}\Biggr{)}$
@@ IV Strong gravitational lensing observables
From this equation, it is evident that when $\psi=\theta_{n}^{0}$, the image aligns with the source. The sign of $\psi$ determines whether the image appears on the same side ($\psi>0$) or the opposite side ($\psi<0$) of the lens. In the scenario where the black hole is nearly aligned with the source and observer ($\psi\approx 0$), and when the observer and lens are positioned equidistantly from the light source ($r_{OS}=r_{S}=2r_{O}$), light deflection occurs in all directions, leading to the formation of Einstein rings [88, 89, 90, 91, 92]. In this case, Eq. (31) simplifies to [93]
@@
To apply the above relations to astrophysical contexts, we now consider the observational data for the supermassive black holes M87* and Sgr A*. Specifically, M87*, with a mass of $(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$, is located at a distance of $r_{O}=16.8\,\mathrm{Mpc}$ from Earth [7, 8], while Sgr A* has a mass of $4^{+1.1}_{-0.6}\times 10^{6}\,M_{\odot}$ and is situated at a distance of $7.97\,\mathrm{kpc}$ from Earth [9, 10]. Using these parameters in Eq. (33), Fig. 4 displays the outermost Einstein rings (corresponding to $n=1$) for both M87* and Sgr A*, assuming that they are PHBHs. These rings are presented in the celestial coordinate system of an observer on Earth, with coordinates $X$ and $Y$, and are computed for different values of the $M/\lambda$ ratio and for specific ranges of the $\mathcal{c}$ parameter.
@@
Figure 4: The outermost Einstein rings for M87* and Sgr A* as PHBHs, for the four cases of the $M/\lambda$ ratio, and different values of the $\mathcal{c}$-parameter corresponding to each of the cases. The solid colored lines correspond to M87*, while the dashed colored lines correspond to Sgr A*. The interior and exterior black dashed circles correspond, respectively, to M87* and Sgr A* in the case that they were SBHs. The unit of length has been chosen as the black hole mass $M$, for each of the black holes.
@@
$M/\lambda$ $\mathcal{c}$ $\theta_{\infty}$($\mu$as) $s$ ($\mu$as) $\theta_{\infty}$($\mu$as) $s$ ($\mu$as) $r_{\mathrm{mag}}$
@@
We observe that, for any fixed $\mathcal{c}$-parameter, an increase in the $M/\lambda$ ratio results in a rise in both the brightness ratio $r_{\mathrm{mag}}$ and the image position $\theta_{\infty}$. On the other hand, when the $M/\lambda$ ratio is fixed, an increase in the $\mathcal{c}$-parameter from negative to positive values leads to a decrease in both $r_{\mathrm{mag}}$ and $\theta_{\infty}$. The only exception occurs for the case of $M/\lambda=1$ and $\mathcal{c}=1.2$. As observed in Fig. 4(a) and Table 1, this case produces smaller Einstein rings compared to other scenarios, and hence, it is considered unreliable within the theory, particularly when observational constraints are applied (see below).
@@
Figure 5: The $\mathcal{c}$-profile of the theoretical shadow angular diameter ($2\theta_{\infty}$) of PHBHs (in $\mu$as), compared with the observed shadow diameter $\theta_{d}$ of M87*, within the $1\sigma$ confidence level, for four different values of the $M/\lambda$ ratio.
@@
Figure 6: The $\mathcal{c}$-profile of the theoretical shadow angular diameter ($2\theta_{\infty}$) of PHBHs (in $\mu$as), compared with the observed shadow diameter $\theta_{d}$ of Sgr A*, within the $1\sigma$ confidence level, for four different values of the $M/\lambda$ ratio.
@@ V Impact of the primary scalar hair on higher-order accretion disk images
As observed, for $n=0$, no half-orbits occur around the black hole, and the light rays do not intersect the equatorial plane (the accretion disk) after their departure. In contrast, for $n\geq 1$, photons perform half-orbits around the black hole, leading to the formation of additional images on the observer’s screen. For $n\geq 2$, as photons complete a full loop, higher-order photon rings are generated, which, as shown in Ref. [53], are known as relativistic images. If the observer is located in the northern hemisphere (i.e., $20\leq\vartheta_{O}\leq\pi/$ 2) and the emission ring of radius $r_{S}$ is assumed to be thin and positioned on the equatorial plane (i.e., $\vartheta_{S}=\pi/2$ and $0\leq\varphi_{S}\leq 2\pi$), the impact parameter in Eq. (37) becomes a function of $\gamma$, which now varies within the range $\pi/2-\vartheta_{O}\leq\gamma\leq\pi/2+\vartheta_{O}$.
@@
which correspond to the energy and angular momentum of the test particles, respectively. Accordingly, the orbit of massive particles is characterized by the Hamilton-Jacobi equation $12\mathcal{L}=-$ 1. The corresponding effective potential is given by
@@ V.1 New constraints on the black hole’s primary scalar hair
The general relativistic magnetohydrodynamics (GRMHD) simulations, based on observational data from the Event Horizon Telescope (EHT) for M87*, reveal a photon ring with an angular diameter of $\theta_{d}=(42\pm 3)\,\mathrm{\mu as}$, located outside the interior shadow [5]. An inclination of $17^{\circ}$ between the approaching jet and the line of sight was adopted, as stated in Ref. [106]. This confirms that the aforementioned ring is distinct from the actual black hole shadow, which is bounded by the photon sphere characterized by $b_{\infty}=b_{c}$. As observed in the diagrams in Fig. 10, for $\vartheta_{O}=17^{\circ}$, the $b_{\mathrm{sec}}$ image appears nearly as a thick ring. Consequently, one can theoretically associate this ring (formed by the $b_{\mathrm{sec}}$ images) with the photon ring, which marks the boundary of the shadow of M87*, assuming it to be a PHBH. It is important to note that in the published image of M87*, this ring is dominated by the direct emission from the accretion disk (i.e., the primary image in Fig. 7) and is therefore not visible. Thus, the generated ring, which borders the shadow, results from analyzing the data based on the GRMHD simulations.
@@
These constraints were derived based on the assumption that the light ring observed in the GRMHD simulation of M87* corresponds not to the boundary of the real inner black hole shadow, but to the primary image of the accretion disk. Moreover, we considered the furthest possible point from the inner shadow by taking the angular position $\varphi=3\pi/2$ on the $b_{\mathrm{sec}}$ ring (see Fig. 10(b,c)), generating conditions similar to those in the GRMHD simulation of M87*. Therefore, the obtained constraints are novel and, from the perspective of the shadow diameter, more reliable. However, referring to the data summarized in Tables 2 and 3, the difference between the proposed constraints is significant, and a consistency between such constraints needs to be established. By comparing these constraints for M87*, one can initially rely on the $M/\lambda=1$ case, as it is the most intuitive and natural choice. Secondly, by comparing the upper and lower limits given in these two tables, it is evident that the most reliable choice is to constrain the primary hair as $-0.2261\leq\mathcal{c}\leq 0.0553$. This domain is covered by all other $M/\lambda$ ratios and is consistent with both approaches used in the observational tests.
It is important to note, however, that as indicated in Ref. [1], PHBHs with their scalar hair parameter within the range $-9.0969\leq\mathcal{c}\leq 1.3191$ are locally unstable. Therefore, the PHBH models proposed in this section, based on the EHT constraints, are also locally unstable.

### v2

@@ I Introduction
Horndeski theory represents the most general class of scalar-tensor theories that maintain second-order field equations while avoiding Ostrogradski instabilities [30]. Later advancements demonstrated that scalar-tensor theories could accommodate higher-order field equations without instability by utilizing degenerate Lagrangians. This led to the development of Beyond Horndeski and Degenerate Higher-Order Scalar–Tensor theories [31, 32, 33, 34, 35, 36, 37, 38, 39, 40], and even rigorous constraints have been proposed [41].
Obtaining black hole solutions in these modified theories is challenging due to extended no-hair theorems, which constrain the existence of non-trivial scalar field configurations [42, 43, 44, 45]. The no-hair conjecture posits that a black hole is entirely described by its mass $M$, angular momentum $J$, and electric charge $Q$, measurable at asymptotic infinity [46]. A black hole with an additional global charge is said to possess primary hair, whereas secondary hair implies that the metric remains fully determined by $M$, $J$, and $Q$ despite the presence of additional fields. Both cases are often categorized under the broader concept of “hairy black holes” [47, 48, 49, 50].
@@
Building on the Virbhadra-Ellis lens equation [56, 57], Bozza extended analytical lensing methods to a wide class of static and spherically symmetric spacetimes, demonstrating that logarithmic divergence of the deflection angle at the photon sphere is a common feature in such backgrounds. He also identified the primary and secondary images and reported the existence of two infinite sets of relativistic images [58]. These findings were later extended by considering finite observer and source positions [59]. In cases where the black hole is illuminated by a thin accretion disk, Tsupko applied Bozza’s method to determine the shape of higher-order images using both analytical and numerical techniques [60]. More recently, Ref. [61] demonstrated that the spacing between higher-order images (termed as photon rings in that reference) can reveal key properties of the underlying spacetime. It is, however, important to clarify the terminology used for the various light rings forming around the black hole shadow. Following Ref. [62], images with order $n\geq 2$ are termed photon rings, while the $n=1$ image is known as the lensing ring or secondary image. In this work, the tertiary image corresponds to the $n=2$ photon ring in that reference. Notably, the general relativistic magnetohydrodynamics (GRMHD) simulations predominantly observe the lensing ring (secondary image), which appears as the photon ring in those simulations. The tertiary image, being thin and closely attached to the inner shadow, remains unresolved in these observations.
@@
Strong gravitational lensing remains a powerful tool for probing black hole spacetimes in both general relativity and alternative theories of gravity. Numerous studies have explored this phenomenon in various contexts [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88].
@@
Motivated by the considerations above regarding the choice of the gravitational theory, we structure this paper as follows: In Sect. II, we provide a concise review of the theory under investigation, introducing the static, spherically symmetric, asymptotically flat black hole solution proposed within this framework and discussing its key properties. In Sect. III, we initiate our analysis by formulating the deflection angle of light in the strong lensing regime, considering both the source and the observer at finite distances from the black hole. The relevant integrals are then evaluated numerically, and we examine how the deflection angle varies with the impact parameter of the passing light rays, as perceived by the observer. Sect. IV is dedicated to the lens equation governing relativistic images of different orders. We utilize this equation to explore the response of RERs to variations in the black hole’s scalar hair. Additionally, we introduce three key observables in strong gravitational lensing and employ the angular diameter of relativistic images as a tool to assess the black hole’s properties in light of the EHT data for M87* and Sgr A*. In Sect. V, we adopt an alternative approach to constraining the black hole parameters by considering a thin accretion disk as the source of illumination. We derive analytical expressions for the shapes of images of different orders and establish a connection between the observed black hole shadow and the primary image of the accretion disk. Using this framework, we impose new constraints on the black hole’s scalar hair. Furthermore, we demonstrate why higher-order images play a less significant role in astrophysical observational tests based on EHT data. Finally, we summarize our findings in Sect. VI.
@@
Throughout this study, we employ geometric units where $G=\hbar=\ell_{p}=c=1$, adopt the sign convention (,+,+,+)(-,+,+,+), and use primes to denote derivatives with respect to the radial coordinate.
@@ II The black hole solution with primary scalar hair
This section outlines the fundamental properties of the static, spherically symmetric black hole solution introduced in Ref. [51]. In the framework of Beyond Horndeski theories [89], where the shift symmetry $\tilde{\chi}\rightarrow\tilde{\chi}+\text{const.}$ and parity symmetry $\tilde{\chi}\rightarrow{-\tilde{\chi}}$ are preserved, the theory is characterized by three arbitrary functions, $G_{2}$, $G_{4}$, and $F_{4}$, which are dependent on the kinetic term of the scalar field, given by $X=-\frac{1}{2}\tilde{\chi}_{,\mu}\tilde{\chi}^{,\mu}$. The corresponding action takes the form
$I\left[g_{\mu\nu},\tilde{\chi}\right]=\frac{1}{16\pi}\int\mathrm{d}^{4}x\,\sqrt{-g}\Biggl{\{}G_{2}(X)+G_{4}(X)R+G_{4X}\left[(\square\tilde{\chi})^{2}-\tilde{\chi}_{\mu\nu}\tilde{\chi}^{\mu\nu}\right]+F_{4}(X)\epsilon^{\mu\nu\rho\sigma}\epsilon^{\alpha\beta\gamma}{}_{\sigma}\tilde{\chi}_{\mu}\tilde{\chi}_{\alpha}\tilde{\chi}_{\nu\beta}\tilde{\chi}_{\rho\gamma}\Biggr{\}},$ (1)
@@
Since the theory is invariant under the transformation $\lambda\rightarrow-\lambda$, we may, without loss of generality, restrict to $\lambda>0$. Setting $M=\mathcal{c}=0$ yields the Minkowski spacetime, while the case $M=0$ with $\mathcal{c}\neq 0$ corresponds to a nontrivial geometry. For $\mathcal{c}<0$, the solution describes a black hole, whereas $\mathcal{c}>0$ – for sufficiently large values, as discussed below – leads to a naked singularity.
@@ III Strong gravitational lensing by the PHBH
where $1\leq\epsilon<\infty$. Since strong lensing effects are significant for $b\approx b_{c}$ (or equivalently, $R\approx r_{p}$), corresponding to trajectories near the photon sphere, the parameter $\epsilon$ is expected to be small. In other words,
@@
Note that, as $R$ approaches $r_{p}$, then $\mathcal{R}$ approaches 0, and hence, the integrals in Eq. (12) diverge. In Refs. [58, 59], a method was developed to address this divergence using logarithmic approximations. Under these approximations, and assuming the source and observer are located at finite positions, the azimuthal angle shift is given by [61]
@@
$\times\Biggl{\{}\sqrt{2}f_{0}\eta\Biggl{[}r_{p}\Biggl{(}-12M\left(\lambda^{2}+r_{p}^{2}\right)^{3}+6r_{p}^{7}+18\lambda^{2}r_{p}^{5}+18\lambda^{4}r_{p}^{3}+6\lambda^{6}r_{p}\Biggr{)}$
@@ IV Strong gravitational lensing observables
From this equation, it is evident that when $\psi=\theta_{n}^{0}$, the image aligns with the source. The sign of $\psi$ determines whether the image appears on the same side ($\psi>0$) or the opposite side ($\psi<0$) of the lens. In the scenario where the black hole is nearly aligned with the source and observer ($\psi\approx 0$), and when the observer and lens are positioned equidistantly from the light source ($r_{OS}=r_{S}=2r_{O}$), light deflection occurs in all directions, leading to the formation of relativistic Einstein rings (RERs) [98, 99, 100, 101, 102, 103]. In this case, Eq. (31) simplifies to [104]
@@
To apply the above relations to astrophysical contexts, we now consider the observational data for the supermassive black holes M87* and Sgr A*. Specifically, M87*, with a mass of $(6.5\pm 0.7)\times 10^{9}\,M_{\odot}$, is located at a distance of $r_{O}=16.8\,\mathrm{Mpc}$ from Earth [7, 8], while Sgr A* has a mass of $4^{+1.1}_{-0.6}\times 10^{6}\,M_{\odot}$ and is situated at a distance of $7.97\,\mathrm{kpc}$ from Earth [9, 10]. Using these parameters in Eq. (33), Fig. 4 displays the outermost RERs (corresponding to $n=1$) for both M87* and Sgr A*, assuming that they are PHBHs. These rings are presented in the celestial coordinate system of an observer on Earth, with coordinates $X$ and $Y$, and are computed for different values of the $M/\lambda$ ratio and for specific ranges of the $\mathcal{c}$ parameter.
@@
Figure 4: The outermost RERs for M87* and Sgr A* as PHBHs, for the four cases of the $M/\lambda$ ratio, and different values of the $\mathcal{c}$-parameter corresponding to each of the cases. The solid colored lines correspond to M87*, while the dashed colored lines correspond to Sgr A*. The interior and exterior black dashed circles correspond, respectively, to M87* and Sgr A* in the case that they were SBHs. The unit of length has been chosen as the black hole mass $M$, for each of the black holes.
@@
[…]
@@
We observe that, for any fixed $\mathcal{c}$-parameter, an increase in the $M/\lambda$ ratio results in a rise in both the brightness ratio $r_{\mathrm{mag}}$ and the image position $\theta_{\infty}$. On the other hand, when the $M/\lambda$ ratio is fixed, an increase in the $\mathcal{c}$-parameter from negative to positive values leads to a decrease in both $r_{\mathrm{mag}}$ and $\theta_{\infty}$. The only exception occurs for the case of $M/\lambda=1$ and $\mathcal{c}=1.2$. As observed in Fig. 4(a) and Table 1, this case produces smaller RERs compared to other scenarios, and hence, it is considered unreliable within the theory, particularly when observational constraints are applied (see below).
@@
[…]
@@
[…]
@@ V Impact of the primary scalar hair on higher-order accretion disk images
As observed, for $n=0$, no half-orbits occur around the black hole, and the light rays do not intersect the equatorial plane (the accretion disk) after their departure. In contrast, for $n\geq 1$, photons perform half-orbits around the black hole, leading to the formation of additional images on the observer’s screen. For $n\geq 2$, as photons complete a full loop, higher-order photon rings are generated, which, as shown in Ref. [59], are known as relativistic images. If the observer is located in the northern hemisphere (i.e., $0\leq\vartheta_{O}\leq\pi/2$) and the emission ring of radius $r_{S}$ is assumed to be thin and positioned on the equatorial plane (i.e., $\vartheta_{S}=\pi/2$ and $0\leq\varphi_{S}\leq 2\pi$), the impact parameter in Eq. (37) becomes a function of $\gamma$, which now varies within the range $\pi/2-\vartheta_{O}\leq\gamma\leq\pi/2+\vartheta_{O}$.
@@
which correspond to the energy and angular momentum of the test particles, respectively. Accordingly, the orbit of massive particles is characterized by the Hamilton-Jacobi equation $2\mathcal{L}=-1$. The corresponding effective potential is given by
@@ V.1 New constraints on the black hole’s primary scalar hair
The GRMHD simulations, based on observational data from the Event Horizon Telescope (EHT) for M87*, reveal a photon ring with an angular diameter of $\theta_{d}=(42\pm 3)\,\mathrm{\mu as}$, located outside the interior shadow [5]. An inclination of $17^{\circ}$ between the approaching jet and the line of sight was adopted, as stated in Ref. [117]. This confirms that the aforementioned ring is distinct from the actual black hole shadow, which is bounded by the photon sphere characterized by $b_{\infty}=b_{c}$. As observed in the diagrams in Fig. 10, for $\vartheta_{O}=17^{\circ}$, the $b_{\mathrm{sec}}$ image appears nearly as a thick ring. Consequently, one can theoretically associate this ring (formed by the $b_{\mathrm{sec}}$ images) with the photon ring, which marks the boundary of the shadow of M87*, assuming it to be a PHBH. It is important to note that in the published image of M87*, this ring is dominated by the direct emission from the accretion disk (i.e., the primary image in Fig. 7) and is therefore not visible. Thus, the generated ring, which borders the shadow, results from analyzing the data based on the GRMHD simulations.
@@
These constraints were derived based on the assumption that the light ring observed in the GRMHD simulation of M87* corresponds not to the boundary of the real inner black hole shadow, but to the primary image of the accretion disk. Moreover, we considered the furthest possible point from the inner shadow by taking the angular position $\varphi=3\pi/2$ on the $b_{\mathrm{sec}}$ ring (see Fig. 10(b,c)), generating conditions similar to those in the GRMHD simulation of M87*. Therefore, the obtained constraints are novel and, from the perspective of the shadow diameter, more reliable. However, referring to the data summarized in Tables 2 and 3, the difference between the proposed constraints is significant, and a consistency between such constraints needs to be established. By comparing these constraints for M87*, one can initially rely on the $M/\lambda=1$ case, as it is the most intuitive and natural choice. Secondly, by comparing the upper and lower limits given in these two tables, it is evident that the most reliable choice is to constrain the primary hair as $0.2261\leq\mathcal{c}\leq 0.0553$. This domain is covered by all other $M/\lambda$ ratios and is consistent with both approaches used in the observational tests.
It is important to note, however, that as indicated in Ref. [1], PHBHs with their scalar hair parameter within the range $9.0969\leq\mathcal{c}\leq 1.3191$ are locally unstable. Therefore, the PHBH models proposed in this section, based on the EHT constraints, are also locally unstable. This local instability highlights the need for considering stabilizing extensions of the theory. In particular, Degenerate Higher-Order Scalar-Tensor (DHOST) frameworks [34, 35] provide promising avenues to embed these solutions within a more general and potentially stable scalar-tensor gravity context. Investigations into the thermodynamic and dynamical stability within such extended theories remain an important direction for future work.